OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x4+ 9x2 f(x) = x2(x + 3i)(x - 3i) f(x) = x2(x + 3i)2 f(x) = x2(3x + i)(3x - i) f(x) = x2(3x + i)2

ganeshie8 (ganeshie8):

factor gcf first

OpenStudy (anonymous):

greatest common facor

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

\(f(x) = x^4+ 9x^2\) \(= x^2(x^2 + 9)\)

OpenStudy (anonymous):

Thats cool didnt know that... whoopsie

ganeshie8 (ganeshie8):

its okay, next find the zeroes of \(x^2+9\)

ganeshie8 (ganeshie8):

and write the zeroes as factors

OpenStudy (anonymous):

x^2 + 9 and x^2 - 9?

ganeshie8 (ganeshie8):

\(x^2 + 9 = 0\) \(x^2 = - 9 \) \(x = \pm 3i \)

ganeshie8 (ganeshie8):

since \(\pm 3i\) are zeroes of \(x^2+9\), \((x+3i)(x-3i)\) are factors of it.

OpenStudy (anonymous):

±3 gotcha

ganeshie8 (ganeshie8):

dont forget \(i\)

OpenStudy (anonymous):

oops

ganeshie8 (ganeshie8):

Overall, \(x^4+ 9x^2 = x^2(x+3i)(x-3i)\)